# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ 4.3. GW Barycenters via Blow-Up: The Synthesis Problem
$\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S}):=\left\{[\mathbb{Y}_{\lambda}]\mid\nabla G_{\lambda}([\mathbb{Y}_{\lambda}])=0\text{ for some }\lambda\in\Delta_{S-1}\right\}.$ (33)
In (32) we consider all minimizers $[\mathbb{Y}_{\lambda}]$ of the general GW synthesis problem (4), with $\lambda\in\Delta_{S-1}$, that is, $[\mathbb{Y}_{\lambda}]\in\operatorname*{arg\,min}_{[\mathbb{Y}]}G_{\lambda}([\mathbb{Y}])$ (also referred to as Fréchet means). The number of nodes in the networks from the set (32), as well as the distribution of mass among them, is not predetermined in advance, unlike in the set $\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s=1}^{S})$ defined in (10). In contrast, the set given by (33) consists of the classes $[\mathbb{Y}_{\lambda}]$ that are critical points of the objective functional in the minimization problem (4), i.e., of the functional $G_{\lambda}$ defined in (31) for some $\lambda\in\Delta_{S-1}$ (also referred to as Karcher means).
@@
  • If $[\mathbb{Y}^{*}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$, then $\mathbb{Y}^{*}\sim^{w}\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$ for some vector $\lambda\in\Delta_{S-1}$.
  • If $[\mathbb{Y}^{*}]\in{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$, then $\mathbb{Y}^{*}\sim^{w}\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$ for some vector $\lambda\in\Delta_{S-1}$.
@@
If $[\mathbb{Y}^{*}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$, then, using Proposition 4.3, we have $0=\nabla G_{\lambda}[\mathbb{Y}^{*}]=\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}-\mathbf{Y}_{b}^{*}$. Thus, $\mathbb{Y}^{*}\sim^{w}(\mathbf{Y}_{b}^{*},\mathrm{q}_{b})=\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$, for some vector $\lambda\in\Delta_{S-1}$.
If $[\mathbb{Y}^{*}]\in{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$ then, due to the first-order optimality condition (see Remark B.10 in App. B.3), $[\mathbb{Y}^{*}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$. Hence, $\mathbb{Y}^{*}\sim^{w}\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$. ∎
In App. B.3, we provide a direct proof of the last bullet point in Lemma 4.5 that does not rely on the first-order optimality condition, but instead builds on the fixed-point approach following the ideas of the proof of Theorem 3.3.
In Lemma 4.5, the blow-up technique can be interpreted as a ‘change of coordinates’, in which all nodes are re-aligned, relabeled, or reordered with respect to those of a GW barycenter $\mathbb{Y}^{*}$. Once this alignment is established, the new templates (which are weakly isomorphic to the original ones) have the same number of nodes and share a common mass distribution $\mathbf{q}_{b}$. The GW barycenter formula (for the weight matrix) will then take the form $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}$, which resembles an Euclidean barycenter171717(see (43) in App. A). This may indicate that the GW framework corresponds to the Euclidean geometry on the space of real square matrices of fixed size ($M_{b}\times M_{b}$). However, there are many caveats. Appropriate mass must be distributed at the nodes of the network $\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$, namely the probability vector $\mathbf{q}_{b}$, which does not only depend on the templates but also on $\mathbb{Y}^{*}$. Also, the order or labels of the coordinates of $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}$ depend on the blow-up template matrices $\{\mathbf{X}_{b}^{s}\}_{s\in[S]}$, which, in turn, are determined by optimal GW assignments with respect to $\mathbb{Y}^{*}$. Moreover, if a new vector $\lambda\in\Delta_{S-1}$ is chosen, the formula of the barycenter does no longer apply and new blow-ups of the templates must be taken respect to a new $\mathbb{Y}^{*}$. For further discussion, see Remark B.13 in App. B.3. We also refer the reader to App. E for an additional discussion on the parallels and differences between Wasserstein and GW barycenters, as well as the methods used to compute them.
@@ 4.4. GW Barycenters via Blow-Up: The Analysis Problem
$\mathcal{A}_{sr}:=\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}^{s}_{b}-\mathbf{Y}_{b})^{T}(\mathbf{X}^{r}_{b}-\mathbf{Y}_{b})\right),\qquad\forall s,r\in[S].$ (34)
@@
This proof is inspired by analysis in [68], and a consequence of Proposition 4.3. Since the networks involved in the statement have finitely many nodes, we can use (26) to express the inner product in the tangent space $\mathrm{Tan}_{[\mathbb{Y}]}$ as a weighted trace of matrices. Indeed, after blow-up all the matrices have the same size $M_{b}\times M_{b}$. Hence, we can consider the variant of Frobenius norm in the space of $M_{b}\times M_{b}$ square matrices weighted by $\mathrm{q}_{b}\otimes\mathrm{q}_{b}$:
@@
$=\sum_{s,r=1}^{S}\lambda_{s}\lambda_{r}\underbrace{\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}^{s}_{b}-\mathbf{Y}_{b})^{T}(\mathbf{X}^{r}_{b}-\mathbf{Y}_{b})\right)}_{\mathcal{\mathcal{missing}}\mathcal{A}_{sr}}$
@@
Let $\{[\mathbb{X}^{s}]\}_{s\in[S]}$ be classes of finite networks. Given a finite class $[\mathbb{Y}]$, consider a blow-up $(\mathbf{Y}_{b},\mathrm{q}_{b})$, $\mathbf{X}^{1}_{b},\dots,\mathbf{X}^{S}_{b}$ as in Algorithm 3, and let $\mathcal{A}$ be the $S\times S$ matrix given by the expression (34) in 4.6. Then $[\mathbb{Y}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s\in[S]})$ if and only if
$\min_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{A}\lambda=0,$ (35)
and the solution argument $\lambda_{\mathbb{Y}}\in\Delta_{S-1}$ is the corresponding weight vector to $[\mathbb{Y}]$ as a weak GW barycenter in the space given in (33). Equivalently, $\lambda_{\mathbb{Y}}$ solves the linear problem
${L\,\lambda=c}\qquad\text{ subject to }\lambda\in\Delta_{S-1}\subset\mathbb{R}^{S},$ (36)
@@
$L_{sr}:=\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}^{s}_{b})^{T}\,\mathbf{X}^{r}_{b}\right)\quad\forall s,r\in[S],\quad c_{s}:=\mathrm{tr}_{\mathrm{q}_{b}}\left(\mathbf{Y}^{T}_{b}\,\mathbf{X}^{s}_{b}\right)\quad\forall s\in[S].$ (37)
@@
First, let us prove that (35) defines a convex quadratic program. On the one hand, the set of restrictions $\Delta_{S-1}$ is a convex set. On the other hand, $\mathcal{A}$ is a symmetric positive semidefinite $S\times S$ matrix because it is the Gram matrix of the system of matrices $\{\mathbf{X}^{s}_{b}-\mathbf{Y}_{b}\}_{s=1}^{S}$ with respect to the inner product in $L^{2}(Y_{b}\times Y_{b},\mu_{Y_{b}}\otimes\mu_{Y_{b}})$ (i.e., $(A,B)\mapsto\mathrm{tr}_{\mathrm{q}_{b}}(A^{T}B)$ for all $A,B\in\mathbb{R}^{M_{b}\times M_{b}}$) [53].
By definition, $[\mathbb{Y}]\in\widetilde{\mathrm{Bary}}(\{(\mathbf{X}^{s})\}_{s=1}^{S})$ if and only if $\nabla G_{\lambda}([\mathbb{Y}])=0$ for some $\lambda\in\Delta_{S-1}$. A vector is the zero vector if and only if its norm is zero, and $\|\nabla G_{\lambda}([\mathbb{Y}])\|_{\mathrm{Tan}_{[\mathbb{Y}]}}=0$ for some $\lambda\in\Delta_{S-1}$ if and only if the problem (35) has a solution $\lambda\in\Delta_{S-1}$ (by 4.6).
@@
where the matrix $L$ and the vector $c$ are given by (37). As a conclusion, the critical points can be expressed as the solutions $\lambda\in\mathbb{R}^{S}$ of the linear system $L\lambda=c$. These solutions might not be unique but under the GW-BCM condition (i.e., assuming $[\mathbb{Y}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$) there exists at least one solution $\lambda_{\mathbb{Y}}$ in $\Delta_{S-1}$. That is, the solutions of (35) are given by (36). ∎
As a direct corollary, under the same hypothesis as Theorem 4.7, if $[\mathbb{Y}]\in{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s\in[S]})$ then (35) is satisfied, and the solution argument $\lambda_{\mathbb{Y}}$ is its corresponding weight vector.
Algorithm 4 implements the analysis scheme laid out in Theorem 4.7, and we refer the reader to App. C for a discussion on its computational complexity.
@@
5:        $\mathcal{\mathcal{missing}}\mathcal{A}_{sr}=\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}^{s}_{b}-\mathbf{Y}_{b})^{T}(\mathbf{X}^{r}_{b}-\mathbf{Y}_{b})\right)$
@@
9:  $\lambda\leftarrow\arg\min_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{A}\lambda$

### v2

@@ 4.3. GW Barycenters via Blow-Up: The Synthesis Problem
(32) $\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S}):=\left\{[\mathbb{Y}_{\lambda}]\mid\nabla G_{\lambda}([\mathbb{Y}_{\lambda}])=0\text{ for some }\lambda\in\Delta_{S-1}\right\}.$
In (31) we consider all minimizers $\mathbb{Y}_{\lambda}]$ of the general GW synthesis problem (2), with $\lambda\in\Delta_{S-1}$, that is, $\mathbb{Y}_{\lambda}]\in\operatorname*{arg\,min}_{[\mathbb{Y}]}G_{\lambda}([\mathbb{Y}])$ (also referred to as Fréchet means). The number of nodes in the networks from the set (31), as well as the distribution of mass among them, is not predetermined in advance, unlike in the set $\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s=1}^{S})$ defined in (7). In contrast, the set given by (32) consists of the classes $\mathbb{Y}_{\lambda}]$ that are critical points of the objective functional in the minimization problem (2), i.e., of the functional $G_{\lambda}$ defined in (30) for some $\lambda\in\Delta_{S-1}$ (also referred to as Karcher means).
@@
  • If $\mathbb{Y}^{*}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$, then $\mathbb{Y}^{*}\sim^{w}\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$ for some vector $\lambda\in\Delta_{S-1}$.
  • If $\mathbb{Y}^{*}]\in{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$, then $\mathbb{Y}^{*}\sim^{w}\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$ for some vector $\lambda\in\Delta_{S-1}$.
@@
If $\mathbb{Y}^{*}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$, then, using Proposition 4.3, we have $0=\nabla G_{\lambda}[\mathbb{Y}^{*}]=\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}-\mathbf{Y}_{b}^{*}$. Thus, $\mathbb{Y}^{*}\sim^{w}(\mathbf{Y}_{b}^{*},\mathrm{q}_{b})=\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$, for some vector $\lambda\in\Delta_{S-1}$.
If $\mathbb{Y}^{*}]\in{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$ then, due to first-order optimality conditions [15] (see Remark B.10 in Suppl. Mat. B.3), $\mathbb{Y}^{*}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$. Hence, $\mathbb{Y}^{*}\sim^{w}\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$. ∎
In Suppl. Mat. B.3, we provide a direct proof of the last bullet point in Lemma 4.5 that does not rely on the first-order optimality condition, but instead builds on the fixed-point approach following the ideas of the proof of Theorem 3.3.
In Lemma 4.5, the blow-up technique can be interpreted as a ‘change of coordinates’, in which all nodes are re-aligned, relabeled, or reordered with respect to those of a GW barycenter $\mathbb{Y}^{*}$. Once this alignment is established, the new templates (which are weakly isomorphic to the original ones) have the same number of nodes and share a common mass distribution $\mathbf{q}_{b}$. The GW barycenter formula (for the weight matrix) will then take the form $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}$, which resembles the classical Euclidean barycenter (see (43) in Suppl. Mat. A). This may indicate that the GW framework corresponds to the Euclidean geometry on the space of real square matrices of fixed size ($M_{b}\times M_{b}$). However, there are many caveats. Appropriate mass must be distributed at the nodes of the network $\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right)$, namely the probability vector $\mathbf{q}_{b}$, which does not only depend on the templates but also on $\mathbb{Y}^{*}$. Also, the order or labels of the coordinates of $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}$ depend on the blow-up template matrices $\{\mathbf{X}_{b}^{s}\}_{s\in[S]}$, which, in turn, are determined by optimal GW assignments with respect to $\mathbb{Y}^{*}$. Moreover, if a new vector $\lambda\in\Delta_{S-1}$ is chosen, the formula of the barycenter does no longer apply and new blow-ups of the templates must be taken respect to a new $\mathbb{Y}^{*}$. For further discussion, see Remark B.13 in Suppl. Mat. B.3. We also refer the reader to Suppl. Mat. F for an additional discussion on the parallels and differences between Wasserstein and GW barycenters, as well as the methods used to compute them.
@@ 4.4. GW Barycenters via Blow-Up: The Analysis Problem
(33) $\mathcal{A}_{sr}:=\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}^{s}_{b}-\mathbf{Y}_{b})^{T}(\mathbf{X}^{r}_{b}-\mathbf{Y}_{b})\right),\qquad\forall s,r\in[S].$
@@
This proof is inspired by the analysis in [72], and a consequence of Proposition 4.3. Since the networks involved in the statement have finitely many nodes, we can use (25) to express the inner product in the tangent space $\mathrm{Tan}_{[\mathbb{Y}]}$ as a weighted trace of matrices. Indeed, after blow-up all the matrices have the same size $M_{b}\times M_{b}$. Hence, we can consider the variant of Frobenius norm in the space of $M_{b}\times M_{b}$ square matrices weighted by $\mathrm{q}_{b}\otimes\mathrm{q}_{b}$:
@@
$=\sum_{s,r=1}^{S}\lambda_{s}\lambda_{r}\underbrace{\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}^{s}_{b}-\mathbf{Y}_{b})^{T}(\mathbf{X}^{r}_{b}-\mathbf{Y}_{b})\right)}_{\mathcal{\mathcal{}}\mathcal{A}_{sr}}$
@@
Let $\{[\mathbb{X}^{s}]\}_{s\in[S]}$ be classes of finite networks. Given a finite class $\mathbb{Y}]$, consider a blow-up $(\mathbf{Y}_{b},\mathrm{q}_{b})$, $\mathbf{X}^{1}_{b},\dots,\mathbf{X}^{S}_{b}$ as in Algorithm 3, and let $\mathcal{A}$ be the $S\times S$ matrix given by the expression (33) in Proposition 4.6. Then $\mathbb{Y}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s\in[S]})$ if and only if
(34) $\min_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{A}\lambda=0,$
and the solution argument $\lambda_{\mathbb{Y}}\in\Delta_{S-1}$ is the corresponding weight vector to $\mathbb{Y}]$ as a weak GW barycenter in the space given in (32). Equivalently, $\lambda_{\mathbb{Y}}$ solves the linear problem
(35) ${L\,\lambda=c}\qquad\text{ subject to }\lambda\in\Delta_{S-1}\subset\mathbb{R}^{S},$
@@
(36) $L_{sr}:=\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}^{s}_{b})^{T}\,\mathbf{X}^{r}_{b}\right)\quad\forall s,r\in[S],\quad c_{s}:=\mathrm{tr}_{\mathrm{q}_{b}}\left(\mathbf{Y}^{T}_{b}\,\mathbf{X}^{s}_{b}\right)\quad\forall s\in[S].$
@@
First, we show that (34) defines a convex quadratic program. On the one hand, the set of restrictions $\Delta_{S-1}$ convex. On the other hand, $\mathcal{A}$ is a symmetric positive semidefinite $S\times S$ matrix as it is the Gram matrix of the system of matrices $\{\mathbf{X}^{s}_{b}-\mathbf{Y}_{b}\}_{s=1}^{S}$ with respect to the inner product in $L^{2}(Y_{b}\times Y_{b},\mu_{Y_{b}}\otimes\mu_{Y_{b}})$ (i.e., $(A,B)\mapsto\mathrm{tr}_{\mathrm{q}_{b}}(A^{T}B)$ $\forall A,B\in\mathbb{R}^{M_{b}\times M_{b}}$) [56].
By definition, $\mathbb{Y}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$ if and only if $\nabla G_{\lambda}([\mathbb{Y}])=0$ for some $\lambda\in\Delta_{S-1}$. A vector is the zero vector if and only if its norm is zero, and $\|\nabla G_{\lambda}([\mathbb{Y}])\|_{\mathrm{Tan}_{[\mathbb{Y}]}}=0$ for some $\lambda\in\Delta_{S-1}$ if and only if the problem (34) has a solution $\lambda\in\Delta_{S-1}$ (by Proposition 4.6).
@@
where the matrix $L$ and the vector $c$ are given by (36). As a conclusion, the critical points can be expressed as the solutions $\lambda\in\mathbb{R}^{S}$ of the linear system $L\lambda=c$. These solutions might not be unique but under the GW-BCM condition (i.e., assuming $\mathbb{Y}]\in\widetilde{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s=1}^{S})$) there exists at least one solution $\lambda_{\mathbb{Y}}$ in $\Delta_{S-1}$. That is, the solutions of (34) are given by (35). ∎
As a direct corollary, under the same hypothesis as Theorem 4.7, if $\mathbb{Y}]\in{\mathrm{Bary}}(\{[\mathbb{X}^{s}]\}_{s\in[S]})$ then (34) is satisfied, and the solution argument $\lambda_{\mathbb{Y}}$ is its corresponding weight vector.
Algorithm 4 implements the analysis scheme laid out in Theorem 4.7, and we refer the reader to Suppl. Mat. C for a discussion on its computational complexity.
@@
5:   $\mathcal{\mathcal{}}\mathcal{A}_{sr}=\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}^{s}_{b}-\mathbf{Y}_{b})^{T}(\mathbf{X}^{r}_{b}-\mathbf{Y}_{b})\right)$
@@
[…]
@@
Remark 4.8.
It is worth mentioning that the blow-up step (Algorithm 3) is essentially equivalent to solving $S$ Gromov-Wasserstein optimization problems. These $S$ concatenated GW problems are defined between networks whose sizes increase as the for-loop progressing from $s=1$ to $s=S$. For simplicity, let us assume that the number of nodes is the same across the templates, i.e, $N^{s}=N$ for all $s\in[S]$. After performing the first blow-up between the first template $\mathbb{X}^{1}$, with $N$ nodes, and $\mathbb{Y}$, with $M$ nodes, we obtain new networks that are weakly isomorphic to the originals but have size $|\mathrm{supp}{(\pi^{1})}|\geq\max\{N,M\}$, where $\pi^{1}\in\Pi(\mathrm{p}^{1},\mathrm{q})$ is an optimal coupling for $GW(\mathbb{X}^{1},\mathbb{Y})$. In the case of classical OT, any optimal coupling has support size at most N+M1N+M-1 (see, e.g., [50, Prop. 3.4]). Linear programming theory does not apply in the GW case due to the quadratic nature of the objective, even though the optimization is performed over the same set of couplings. Best-case scenarios for GW include situations in which an optimal plan is a permutation matrix (with support size $N$, when $N=M$), or, more generally, lies at an extreme point of the transport polytope $\Pi(\mathrm{p}^{1},\mathrm{q})$ and thus has at most N+M1N+M-1 nonzero entries. However, no support-size bound analogous to that in classical OT is currently known for GW minimizers, other than the trivial bound $NM$. For empirical insight, we refer the reader, for e.g., to [15, Suppl. Mat., Sec. C; Figs. 9 &\& 10], where the authors observe that representations often occur in a space much smaller than the naive requirement of size $NM$, in fact of order $N+M$. As a conclusion, as this process continues similarly for subsequent steps, involving blow-ups for the remaining templates $\mathbb{X}^{2},\dots,\mathbb{X}^{S}$ with respect to the updated networks $\mathbb{Y}$, the number of nodes at each iteration $s\in[S]$ increases as well: In the worst case –when the GW plans are dense– the size of the barycenter can be as large as $(NM)^{S}$. Even for sparse plans, an exponential growth of order $(M+N)^{S}$ is, in principle, possible. However, experimental evidence [15] supports the ‘linear’ hypothesis that $M_{b}\approx SN+M$, and this has been corroborated in our experiments (cf. Section 5).